The area of a rectangle with a side length of 24 and the diagonal of 25 is
Answer by stanbon(75887)

The area of a rectangle with a side length of 24 and the diagonal of 25 is
----
other side = sqrt[25^2-24^2] = sqrt(49) = 7
--------------------
Area = 7*24 = 148 sq units
----------------------
Cheers,
Stan H.
========================